OpenStudy (anonymous):

Use the given conditions to write an equation for the line in point-slope form. Passing through (-5, -7) and (-8, -6) A. y - 7 = -1/3 (x - 5) or y - 6 = - (x - 8) C. y + 7 = -1/3 (x + 5) or y + 6 = - (x + 8) D. y + 7 = - 1/3 x - 5 or y + 6 = - x + 7

OpenStudy (anonymous):

@AnswerMyQuestions @Andertj21 @August25 @bless

OpenStudy (texaschic101):

first we find the slope by using the slope formula slope(m) = (y2 - y1) / (x2 - x1) (-5,-7)...x1 = -5 and y1 = -7 (-8,-6)..x2 = -8 and y2 = -6 now we sub slope(m) = (-6 - (-7) / (-8 - (-5) slope(m) = (-6 + 7) / (-8 + 5) slope(m) = 1/-3 slope(m) = -1/3 point slope form : y - y1 = m(x - x1) slope(m) = -1/3 (-8,-6)..x1 = -8 and y1 = -6 now we sub y - (-6) = -1/3(x - (-8) y + 6 = -1/3(x + 8) OR y - y1 = m(x - x1) slope(m) = -1/3 (-5,-7)...x1 = -5 and y1 = -7 now we sub y - (-7) = -1/3(x - (-5) y + 7 = -1/3(x + 5)

OpenStudy (texaschic101):

I would go with C, except the slope is missing in part of the answer

OpenStudy (anonymous):

Thank you so much @texaschic101 you are a big help! :)

OpenStudy (texaschic101):

I try :)

OpenStudy (anonymous):

;) I need help

OpenStudy (anonymous):

@roman12334 can you please post your question in your post Thanks